# A GREAT Architecture for Edge-Based Graph Problems Like TSP

### v2

@@ 3.2 Mathematical Formulations
$x_{i}^{\ell}=\sum_{j\in N(i)}(\alpha^{\prime}_{i,j}W^{\prime}_{1}e_{i,j}^{\ell-1}||\alpha^{\prime\prime}_{i,j}W^{\prime\prime}_{1}e_{j,i}^{\ell-1}),\quad\text{where}$
@@
We note that $W^{\prime}_{1},W^{\prime\prime}_{1},W^{\prime}_{2},W^{\prime\prime}_{2},W^{\prime}_{3},W^{\prime\prime}_{3},W^{\prime}_{4},W^{\prime\prime}_{4}$ are trainable weight matrices of suitable dimension, $d$ is the hidden dimension and ||||| | denotes concatenation. $W^{\prime}_{1}e_{i,j}^{\ell-1},W^{\prime}_{2}e_{i,j}^{\ell-1}$ and $W^{\prime}_{3}e_{i,j}^{\ell-1}$ correspond to the “values”, “keys” and “queries” of the original transformer architecture.
@@
Note that the edge feature consists of four individual terms that are concatenated. Due to the attention mechanism, these terms summarize information on all edges outgoing from node $i$, ingoing to node $i$, outgoing from node $j$, and ingoing to node $j$. The distinction between in- and outgoing edges is again necessary due to asymmetric graphs. The $\alpha^{\prime}$ and $\alpha^{\prime\prime}$ scores are computed as for the node-based GREAT variant.
Feedforward (FF) Subayer: Like in the original transformer architecture, the FF layer has the following form.
@@
where $W_{1},b_{1},W_{2},b_{2}$ are trainable weight matrices and biases of suitable sizes. Moreover, again like in Vaswani et al. (2017), the feedforward sublayers have internal up-projections, which temporarily double the hidden dimension before scaling it down to the original size.
@@
3.3 GREAT-based RL Framework
For our experiments (see Section 4), we built end-to-end RL frameworks for the TSP, CVRP and OP. We train a GREAT-based model to incrementally build solution tours by adding one node at a time to a partial solution. Our framework follows the encoder-decoder approach (where GREAT serves as the encoder and a multi-pointer network as the decoder) and is trained using POMO (Kwon et al. 2020). We visualize the framework in Figure 5. Given an input graph of edge features, GREAT learns edge encodings that reflect the importance of individual edges. Next, the edge encodings are transformed into node encodings by aggregation. While it is possible to create edge-based decoder architectures, converting the edge encodings into node encodings allows us to freely plug in GREAT into existing training pipelines. Further, it reduces the memory requirement from $\mathcal{O}(n^{2})$ edges to $\mathcal{O}(n)$ nodes during decoding. We visualize in Figure 5 also how GREAT potentially learns: We imagine that the edge encodings produced by GREAT reflect how promising edges are to be part of a TSP solution. In the visualization, we assign darker colors to the encodings of such promising edges. Then, when transformed into node encodings, we assume that the edge information gets passed on and node encodings reflect which nodes are connected by important edges. For more details, we refer to the SM.
@@ 3.3 GREAT-based RL Framework
We note that we can apply GREAT to extensions of the TSP such as the CVRP or OP even though capacities and prizes are node-level features. This is because we can easily transform them into edge features. Consider CVRP where a node $j$ has a demand $c_{j}$. We can simply add demand $c_{j}$ to all edges $e_{i,j}$ in the graph. This is because we know that if we have an edge $e_{i,j}$ in the tour, we will visit node $j$ in the next step and therefore need a free capacity in our vehicle big enough to serve the demand of node $j$ which is $c_{j}$. An analogous extension works for OP.
4 Experiments: Learning to Solve Non-Euclidean Routing Problems
@@
While many other studies focus on developing frameworks to solve larger and larger Euclidean routing problem instances, we mostly focus on instances of “only” $100$ nodes in much more challenging distributions in our experiments. However, we also fine-tuned our trained TSP models on TSP instances up to $500$ nodes in a few-shot CL setting to be applicable to larger instances (up to $1,000$ nodes). In detail, we train models for TSP, CVRP and OP on the following “distance”-distributions:
@@
    Asymmetric distribution with triangle inequality (TMAT) as used in (Kwon et al. 2021). We use the code of (Kwon et al. 2021) to generate instances. However, we normalize the distance matrices differently: Instead of a fixed scaling value, we normalize each instance individually such that the biggest distance is exactly 1. By this, we ensure that the distances use the full range of the interval (0,1) as well as possible.
@@
    Extremely asymmetric TSP (XASY) where all pairwise distances are sampled uniformly at random from the interval (0,1). The same distribution was used in (Gaile et al. 2022). Here, the triangle inequality does not hold.
@@
The exact setting for this experiment is the following. For each distribution, we train a GREAT NF and a GREAT NB model with hidden dimension 128. All networks have 5 hidden layers and 8 attention heads. Training is done for 400 epochs and there are 25,000 instances in each epoch. Every 10 epochs, we change the dataset to a fresh set of 25,000 instances (meaning $400\times 25,000=10,000,000$ instances in total). We evaluate the model after each epoch on a distinct validation dataset of $1,000$ instances and save the model with the best validation loss during these 400 epochs for testing. Furthermore, while training, the distances of all instances in the current data batch were multiplied by a factor in the range (0.5, 1.5) to ensure the models learn from a more robust data distribution. This allows us to augment the dataset at inference, as was done in other studies such as Kwon et al. (2020). We note that this data augmentation makes the model learn “out of distribution”, much more than the augmentation method used in e.g., Kwon et al. (2020) which works by rotating Euclidean coordinates. This offers the advantage of potentially increasing the generalization performance of our model at the cost of leading to lower performance gains through augmentation at inference in the original distribution.
@@
The overall framework to construct solutions, as well as the decoder to decode the encodings provided by GREAT and the loss formulation, are adapted from Jin et al. (2023b). We also retrain MatNet (Kwon et al. 2021) and Pointerformer (Jin et al. 2023b) (EUC only) with the exact same hyperparameters and train dataset as GREAT, to create fair benchmarks.
@@
[…]
@@
Table 1: Benchmarking GREAT against neural methods and traditional baselines

### v3

@@ III-B Mathematical Formulations
$x_{i}^{\ell}=\sum_{j\in N(i)}(\alpha^{\prime}_{i,j}W^{\prime}_{1}e_{i,j}^{\ell-1}||\alpha^{\prime\prime}_{i,j}W^{\prime\prime}_{1}e_{j,i}^{\ell-1})$
where
@@
$=\frac{\exp\big(\frac{(W^{\prime}_{2}e_{i,j}^{\ell-1})^{\top}W^{\prime}_{3}e_{i,j}^{\ell-1}}{\sqrt{d}}\big)}{\sum_{k\in N(i)}\exp\big(\frac{(W^{\prime}_{2}e_{i,k}^{\ell-1})^{\top}W^{\prime}_{3}e_{i,k}^{\ell-1}}{\sqrt{d}}\big)}$
@@
$=\frac{\exp\big(\frac{(W^{\prime\prime}_{2}e_{j,i}^{\ell-1})^{\top}W^{\prime\prime}_{3}e_{j,i}^{\ell-1}}{\sqrt{d}}\big)}{\sum_{k\in N(i)}\exp\big(\frac{(W^{\prime\prime}_{2}e_{k,i}^{\ell-1})^{\top}W^{\prime\prime}_{3}e_{k,i}^{\ell-1}}{\sqrt{d}}\big)}$
@@
We note that $W^{\prime}_{1},W^{\prime\prime}_{1},W^{\prime}_{2},W^{\prime\prime}_{2},W^{\prime}_{3},W^{\prime\prime}_{3},W^{\prime}_{4},W^{\prime\prime}_{4}$ are trainable weight matrices of suitable dimension, $d$ is the hidden dimension and $||$ denotes concatenation. $W^{\prime}_{1}e_{i,j}^{\ell-1},W^{\prime}_{2}e_{i,j}^{\ell-1}$ and $W^{\prime}_{3}e_{i,j}^{\ell-1}$ are similar to the “values”, “keys” and “queries” of the original transformer architecture.
@@
Note that the edge feature consists of four individual terms that are concatenated. Due to the attention mechanism, these terms summarize information on all edges outgoing from node $i$, ingoing to node $i$, outgoing from node $j$, and ingoing to node $j$. The distinction between in- and outgoing edges is again necessary due to asymmetric graphs. The $\alpha^{\prime}$ and $\alpha^{\prime\prime}$ scores are computed as for the node-based GREAT variant above.
Remark that the attention used in GREAT (NB and NF) could also be interpreted as a form of gating and it is more similar to the type of “attention” used in architectures like the attention free transformer proposed by [39] that do not use self-attention. Computing real self-attention scores (as used in the original transformer paper [5]) would require us to compare each of the $\mathcal{O}(n^{2})$ edges in a routing problem graph to its $\mathcal{O}(n)$ adjacent edges. This would result in calculating $\mathcal{O}(n^{3})$ many attention scores (compared to $\mathcal{O}(n^{2})$ the way we do it) and lead to a considerable computational increase.
Feedforward (FF) Sublayer: Like in the original transformer architecture, the FF layer has the following form.
@@
where $W_{1},b_{1},W_{2},b_{2}$ are trainable weight matrices and biases of suitable sizes. Moreover, again like in [5], the feedforward sublayers have internal up-projections, which temporarily double the hidden dimension before scaling it down to the original size.
@@
Edge Feature Inputs GREAT Sublayer Add & Norm Feed Forward Add & Norm Edge EmbeddingsGREAT Layer $\times n$
Figure 6: A GREAT layer with sublayers and normalizations
@@
IV Experiments: Learning to Solve Non-Euclidean Routing Problems
IV-A GREAT-based RL Framework
For our experiments, we built end-to-end RL frameworks for the TSP, CVRP and OP. Our framework follows the encoder-decoder approach (where GREAT serves as the encoder and the multi-pointer network adapted from [8] as the decoder). We train our models to incrementally build solution tours by adding one node at a time to a partial solution.
@@ IV-A GREAT-based RL Framework
We point out that we can apply GREAT (which is purely edge-based) to extensions of the TSP such as the CVRP or OP even though capacities and prizes are node-level features. This is because we can easily transform them into edge features. Consider CVRP where a node $j$ has a demand $c_{j}$. We can simply add demand $c_{j}$ to all edges $e_{i,j}$ in the graph. This is because we know that if we have an edge $e_{i,j}$ in the tour, we will visit node $j$ in the next step and therefore need a free capacity in our vehicle big enough to serve the demand of node $j$ which is $c_{j}$. An analogous extension works for OP.
IV-A1 Intuition of the Framework
We visualize the framework in Figure 7 and also show how GREAT potentially learns and passes on information. For the sake of simplicity, we illustrate the idea of the framework for an Euclidean TSP instance. Non-Euclidean instances (and other routing problems) can be processed in the same way. Given an input graph of edge features (e.g., the distances between TSP nodes), GREAT learns edge encodings that reflect the importance of individual edges (i.e., how promising they are to be part of the TSP solution). In the visualization, we assign darker colors to the encodings of such promising edges. Next, the edge encodings are transformed into node encodings by aggregation. This is simply achieved by using a NB GREAT layer that returns the (otherwise temporary) intermediate internal node encodings (note that only the very last GREAT layer returns these node embeddings, otherwise, only edge encodings are passed on to the next layer). Since NF GREAT does not have temporary node embeddings, we use a final NB GREAT layer in our RL framework even when the other GREAT layers are node-free. During this transformation, we assume that the edge information gets passed on and node encodings reflect which nodes are connected by important edges. Our assumption that nodes connected by important edges have similar encodings is backed up by the heatmap visualization in Figure 8. The visualization shows Euclidean distances and cosine similarities between the vector encodings for the nodes in the TSP instance that are returned by the GREAT encoder to be passed on to the decoder. Red frames around a tile in the heatmaps signal that there is an edge between the two nodes in the optimal TSP solution. We can see that for the cosine similarity, the red frames are mostly around tiles with high cosine similarity. Analogously, for the Euclidean distance heatmaps, we can see that the red frames are mostly around tiles with low distances. For generating these heatmaps, we use a node-based GREAT encoder for Euclidean TSP with 100 nodes trained for our experiments. Even though the model is trained on TSP instances of size 100, we can see that the heatmaps indicate similar patterns for TSP instances of size 50 and 30. In Figure 7, we assign similar colors to nodes that are connected by edges deemed promising by GREAT to visualize the effect shown in the heatmaps. In the last step of our encoder-decoder framework, we hypothesize that the decoder can construct solutions from these embeddings by iteratively selecting similar node encodings.
Although it would have been possible to create an edge-based decoder instead of transforming the edge-features produced by the GREAT encoder into node features, converting the edge encodings into node encodings allows us to freely plug in GREAT into existing training pipelines. Further, it reduces the memory requirement from $\mathcal{O}(n^{2})$ edges to $\mathcal{O}(n)$ nodes during decoding.
IV-A2 RL Environment
The rewards for a solution in the environments of the RL framework correspond to the negative traveled distance for TSP and CVRP (ensuring we minimize distance) and to the collected prizes in OP. The rewards of all routing problems are used to formulate losses in a POMO-based setting [7], using multiple rollouts as a robust baseline.
To ensure valid solutions to all considered routing problems, we use masking operations. This means we prevent the model to visit nodes multiple times (TSP, OP and CVRP with exception of the depot), to respect the vehicle capacity (CVRP) and the maximum travel distance (OP). A special case here is the OP in combination with the earlier mentioned XASY distribution. There, it is not simply possible to mask invalid moves that lead to a violation of the maximum travel distance of the OP instance. Since the triangle inequality does not hold for XASY, we cannot rely on simple additive checks to ensure that returning to the depot is still possible. Consider a partial OP solution where we would like to visit a further customer $j$ starting from our current position at node $i$. Even if the direct distances from $i$ to $j$ and from $j$ to the depot appears infeasible (i.e., the sum of these distances would violate the still available distance budget), the actual shortest return path might respect the constraint. Consequently, we would have to compute the true shortest return path to the depot from each candidate node $j$ in order to check feasibility. This would require solving a series of shortest path problems during the decoding process (since after each selection step the shortest path might change as we cannot visit nodes multiple times), which would result in an enormous computational overhead. Therefore, for OP on the XASY distribution, we do not mask moves that could potentially violate maximum travel distance and we require the model to learn valid moves itself. We do this by setting the reward to 0 in case of a violation, incentivizing the model to not take such moves since the goal is to maximize reward.
(a) Cosine similarity TSP30
(b) Cosine similarity TSP50
(c) Cosine similarity TSP100
(d) Euclidean distance TSP30
(e) Euclidean distance TSP50
(f) Euclidean distance TSP100
Figure 8: Vector similarities for node encodings returned by the GREAT encoder in the RL framework for TSP
IV-A3 Experimental Setup
@@
While many other studies focus on developing frameworks to solve larger and larger Euclidean routing problem instances, we mostly focus on instances of “only” $100$ nodes in much more challenging distributions in our experiments. In detail, we train models for TSP, CVRP and OP on the following “distance”-distributions:
@@
    Asymmetric distribution with triangle inequality (TMAT) as used in [9].
@@
    Extremely asymmetric TSP (XASY) where all pairwise distances are sampled uniformly at random from the interval (0,1).
  4. 4.
    A mixed distribution which consists of $\frac{1}{3}$ EUC, TMAT and XASY, respectively, to obtain trained models applicable to all distributions.
@@
The exact setting for this experiment is the following. For each distribution, we train a GREAT NF and a GREAT NB model with hidden dimension 128. All networks use 5 hidden layers and 8 attention heads. We consider these sizes a reasonable trade-off between expressiveness and resource demands, since larger networks are more powerful but more costly to train and run. We train for for 400 epochs and there are 25,000 instances in each epoch. Every 10 epochs, we change the dataset to a fresh set of 25,000 instances (meaning $400\times 25,000=10,000,000$ instance presentations in total). We used the ADAM optimizer [40] with a learning rate of $0.0001$ to update the network parameters (we found this learning rate to empirically work best in our setting). After each epoch, we evaluate the model on a distinct validation dataset of $1,000$ instances and save the model with the best validation loss during the 400 epochs of training. Furthermore, while training, the distances of all instances in the current data batch are multiplied by a factor in the range (0.5, 1.5) to ensure learning from a more robust data distribution. This allows us to augment the dataset at inference by multiplying the original distances with several factors and solving the resulting scaled instances multiple times. A similar idea was followed in other studies such as [7]. We note that our form of data augmentation by scaling makes the model learn “out of distribution”, much more than the augmentation method used in e.g., [7] which works by rotating Euclidean coordinates. This offers the advantage of potentially increasing the generalization performance of our model as it has encountered distance matrices of very different scales during training. However, this comes at the cost of leading to lower performance gains through augmentation at inference in the original distribution (since the “focus” of the model will by on instances scaled with a factor close to 1, i.e., the original, unscaled instances) compared to the augmentation of [7].
@@
The overall framework to construct solutions, as well as the decoder to decode the encodings provided by GREAT and the loss formulation, are adapted from [8].
We also retrain MatNet [9] and Pointerformer [8] (EUC only) with the exact same hyperparameters and train dataset as GREAT, to create fair benchmarks and comparisons.
IV-B Curriculum Learning
While the main experimental analysis of our work focuses on creating competitive models for TSP, CVRP and OP with 100 nodes on challenging, asymmetric distributions, we also explore tackling larger TSP instances (up to $1000$ nodes) by using a few-shot CL set-up. This means we fine-tune the trained models of TSP for 100 nodes (TSP100) by retraining for one epoch each on the following instance sizes (10% increment steps from the previous instance size and additionally 200 + 500): [110, 121, 133, 146, 161, 177, 194, 200, 214, 235, 259, 285, 313, 345, 379, 417, 459, 500]. For each instance size, we create a dataset of $2,000$ instances. For instance size 200 and 500 we trained for not only one but five epochs (using the same dataset of $2,000$ in each epoch) and the fine-tuned model was saved afterwards. We note that $2,000$ instance presentations is $0.02\%$ of the $10,000,000$ instance presentations that were used to train the “base models” for TSP100, making the fine-tuning “few-shot”. The datasets were used to train the models in the same end-to-end RL framework used for the experiments with 100 nodes. Since the memory requirement and runtime of such end-to-end RL frameworks is quadratic in the number of nodes, it is essential that our model can be fine-tuned using only such small datasets.
While fine-tuning, we kept the POMO size (i.e., the number of samples trajectories, compare [7] for details) to 100 (due to memory constraints), while setting it to the actual instance size during evaluation. We set the batch size to 16 for instance size <250<250, to 8 for instance size <350<350 and 4 otherwise to respect the available VRAM of our used GPU (a single NVIDIA A40 with 48GB of VRAM).
IV-C Training times & Parameters
In the following, we report the training times for training the CVRP models with 100 nodes (other problems might have slightly different times) in Table I. We trained our models for 400 epochs. The total times also includes validating the models after each epoch (on 1000 instances using x8/x9 augmentation) to be able to save the best-performing model after any epoch as our final model and the times for dataset generation (a new train dataset is used every 10 epochs). Therefore, the total time equals Epoch400+Val400+40DataGenEpoch*400+Val*400+40*DataGen where DataGen is approximately 6 minutes (independent of the used model). All training and inference of neural models has been performed on NVIDIA A40 GPUs with 48GB of VRAM. Additionally to training times, we also provide the number of parameters of each model in Table I.
TABLE I: Training times of the considered architectures
Training Times Epoch Val Total Params
GREAT NB $\simeq$ 7.5min $\simeq$ 60s $\simeq$ 60h 1.42M
GREAT NF $\simeq$ 6.9min $\simeq$ 50s $\simeq$ 55h 1.29M
Pointerformer $\simeq$ 2.1min $\simeq$ 12s $\simeq$ 19h 1.09M
MatNet $\simeq$ 3.4min $\simeq$ 22s $\simeq$ 29h 1.42M
In addition to benchmarking GREAT against the retrained MatNet and Pointerformer networks, we also provide several non-learning-based baselines in our experiments. All such baselines were executed on an Apple M4 10-core CPU (at the time of the experiments among the fastest in consumer devices [41]). However, in case of the EA4OP heuristic for the OP, only 1 instance could be solved at a time due the the program running in a docker environment.
TABLE II: Benchmarking GREAT against neural methods and traditional baselines
@@ IV-C Training times & Parameters
Nearest Neighbor 24.86% 1.3s 36.04% 1.2s 185.26% 1.2s
Nearest Insertion 21.8% 27s 31.8% 28s 301.65% 28s
Farthest Insertion 7.66% 32s 23.92% 33s 310.98% 33s
BQ-NCO $\textbf{0.01\%}^{*}$ $\text{32m}^{*}$ $\textbf{0.96\%}^{*}$ $\text{19m}^{*}$ UNK UNK
@@
GREAT MIX NF x1 4.45% 64s 5.77% 64s 45.78% 64s
GREAT MIX NF x9 3.5% 563s 4.63% 554s 33.32% 552s
GREAT MIX NF x33 3.2% 2017s 4.39% 2014s 29.8% 2015s
GREAT MIX NB x1 2.53% 70s 4.99% 70s 45.97% 70s
GREAT MIX NB x9 1.9% 612s 3.85% 608s 33.19% 609s
GREAT MIX NB x33 1.84% 2212s 3.7% 2211s 30.74% 2207s
@@
GREAT MIX NF x1 4.12% 64s 6.37% 64s 14.83% 64s
GREAT MIX NF x9 3.38% 543s 5.11% 536s 8.04% 542s
GREAT MIX NF x33 3.15% 1961s 4.8% 1956s 5.42% 1962s
GREAT MIX NB x1 3.98% 70s 5.49% 71s 11.88% 71s
GREAT MIX NB x9 3.18% 594s 4.35% 598s 4.91% 604s
GREAT MIX NB x33 3.03% 2148s 4.15% 2183s 3.18% 2199s
@@
GREAT MIX NF x1 -16.97% 58s -24.43% 60s -15.84% 55s
GREAT MIX NF x9 -18.59% 493s -26.07% 498s -21.06% 478s
GREAT MIX NF x33 -19.09% 1774s -26.38% 1806s -21.78% 1742s
GREAT MIX NB x1 -18.22% 65s -25.5% 66s -19.34% 62s
GREAT MIX NB x9 -19.56% 548s -27.02% 556s -23.7% 537s
GREAT MIX NB x33 -19.96% 1995s -27.35% 2032s -24.08% 1928s
@@
TABLE III: Generalization results on test datasets with 30 instances from [34]